1.The math grades on the final exam varied greatly. Using the scores below, how many scores were within one standard deviation of the mean? How many scores were within two standard deviations of the mean?
99 34 86 57 73 85 91 93 46 96 88 79 68 85 89

Answer by Edwin McCravy(20066)

Go to this site:

http://www.calculator.net/standard-deviation-calculator.html

Type in your values with commas between them like this:
99,34,86,57,73,85,91,93,46,96,88,79,68,85,89

Click on "Calculate":

Read these two values:

Mean (Average):	77.933333333333
Population Standard Deviation, 18.382842241854

how many scores were within one standard deviation of the mean?
Add 1 times the standard deviation 18.382842241854 to the mean

77.933333333333 + 1x18.382842241854 = 96.316175575187

Subtract 1 times the standard deviation 18.382842241854 from the mean

77.933333333333 - 1x18.382842241854 = 59.550491091479

List the scores that are between those two numbers:

86,73,85,91,93,96,88,79,68,85,89

Count them:

Answer 11.

------------------------------

how many scores were within two standard deviations of the mean?
Add 2 times the standard deviation 18.382842241854 to the mean

77.933333333333 + 2x18.382842241854 = (you do it)

Subtract 2 times the standard deviation 18.382842241854 from the mean

77.933333333333 - 2x18.382842241854 = (you do it)

List the scores between those two numbers:

(you do it)

Count them:

(you do it).

Edwin